you have just taken out a 30 year $120,000 mortgage on your new home. This mortgage is to be repaid in 360 equal monthly installments. if the stated (nominal) annual interest rate is 15.73 percent, what is the amount of interest portion of the FIRST installment?
Answer #1

The equal monthly installment on the loan is calculated as follows

PMT = Present value of the loan 1-(1 + interest rate) -nt interest rate

$120000 PMT = - 1-(1+0.1573) -360 0.1572 12

Monthly payment on the loan = $1,587.61

Amount of interest portion of the first installment = $ 1573

Month Beginning balance Monthly payment Interest Principal Ending balance
1 120000 1587.61 (15.73 % \div 12) \times 120000 = 1573 1587.61 - 1573 = 14.61 120000 - 14.61 = 119985.39

Month Beginning balance Monthly payment Interest Principal Ending balance 120000 1587.61 1573 14.61 119985.39